For about what value(s) of x does g(x) = k(x) if g(x) = 2x^2 -8 and
k(x) = -x^2 -2?
a. x = 2,-2
b. x = 2.3, -2.3
c. x = 1.4
d. x = 1.4, -1.4

Answer by nabla(475)

The question is where is:
2x^2 -8=-x^2 -2
Solve:
3x^2-6=0
x^2-2=0
x must be +/- sqrt(2)
d would be the approximate answer.